Hi everyone, could you help me solve this equation?
Thank you so much,
Alexa
26. The area of a playground is 221 yd^2. The width of the playground is 4 yd longer than its length. Find the length and width of the playground.
(Hint: Include correct units in your final answer.) (3 points) Answer by ankor@dixie-net.com(22740) 
 26. The area of a playground is 221 yd^2. The width of the playground is 4 yd longer than its length.
Find the length and width of the playground.
:
L*W = 221
It says,"The width of the playground is 4 yd longer than its length." therefore
W = (L+4)
so we can replace W:
L(L+4) = 221
L^2 + 4L = 221
L^2 + 4L - 221 = 0
You can solve this using the quadratic formula, but it will factor to:
(L+17)(L-13) = 0
the positive solution
L = 13 yds is the length
then
13 + 4 = 17 yds is the width
:
:
Check by finding the area with these dimensions:
13 * 17 = 221 sq/yds
:
Sorry about the wrong units here